The length of a rectangel is 3 cm more than 2 times its width. If the area of the rectangle is 99 cm^2, find the dimensions of the rectangle to the nearest hundredth.
Thanks

Answer by checkley71(8403)

(2X+3)X=99 OR 2X~2+3X-99=0 OR USING THE QUADRATIC EQUATION WE GET X=6.33 & 2X+3=15.66 SO 6.33*15.66=99 OR 99~99